# F-91191 Gif-sur-Yvette cedex, France

### v1

@@ 4.3.3 The NP scalars: analytic contributions
In particular, the integral over the azimuthal angle vanishes unless the integrand is independent of $\phi$. Since the angle $\phi$ enters only via the complex combination of coordinates $x\pm iy=e^{\pm i\phi}\sqrt{1-z^{2}}$ and, as discussed, the coefficient functions $\alpha_{i}$ are regular at $z=1$, the expansion of the NP scalar $\Psi_{4}$ around $z=1$ is
@@
This concludes the main part of our proof that the part of the momentum-space metric that is analytic in $\omega$ yields a space-time metric that obeys Sachs’s peeling property.
@@
Each of the coefficients $B$ and $C$ is at least ${\cal O}(G/|\vec{x}|)$ because the first correction to the metric, corresponding to the direct sum of two noninteracting Schwarzschild black holes, is of this order. Moreover, we can use the leading values of the other two coefficients, $A=D=1$. Thus, while the falloff of $\Psi_{i\geq 1}$ is slower than the expected falloff of $\Psi_{0}$, the fall-off of the coefficients $B$ and $C$ makes up for the difference. Consequently, the conclusion of Sec. 4.3.3 that the analytic contributions to the metric exhibit the peeling property are unaffected by redefinitions of the tetrad. We expect that this holds beyond the leading order in the deformation, so that such mixing of NP scalars due to the nontriviallity of the nullness and orthogonality constraints on the tetrad does not, on its own, lead to departures from the peeling property of the Weyl tensor.
In this section we focused on the part of the momentum-space metric that is analytic in $k^{2}$. Thus, the only singularity in this part of the metric originates in the outgoing-graviton propagator. Together with the second essential ingredient, the $z\rightarrow 1$ behavior of the coefficient functions $\alpha_{i}$ in Eq. (46), it led to the position-space metric obeying Sachs’s peeling property. In the next section, we will discuss nonanalytic features of the final-state graviton one-point function and their consequences on the large-distance fall-off of the NP scalars.
@@ 4.4 Radiation region: non-analytic contributions and new peeling violation
To begin, let us recall that since the integrands of final-state graviton one-point function and of the corresponding scattering amplitudes are rational functions. The only difference between them are proportional to integer powers of the squared graviton momentum, $k^{2}$, or to $\varepsilon_{\mu\nu}(k)k^{\nu}$. We chose the tensor factor in Eqs. (22) and (23) so that it is proportional to $k^{2}$ when contracted with the graviton momentum. Moreover, any terms linear in $k^{\mu}$ that survive after contraction drop out when evaluating the Weyl tensor and the NP scalars. Thus, the difference between the contribution of the integrands of the amplitude and graviton one-point function to Eqs. (22) and (23) is effectively an analytic function of $k^{2}$. While, upon integration, these terms may contain non-analytic parts, the $k^{2}$ multiplicative factor guarantees that they cannot contribute to the same power of $1/|\vec{x}|$ as the non-analytic terms that arise from integrating the on-shell integrand without imposing on-shellness for the outgoing graviton. The relevant terms were constructed explicitly in Sec. 2.2 at ${\cal O}(G^{3})$, see Eq. (27). The non-analytic terms were also connected to all orders in Newton’s constant to the dimensional regulator dependence of Weinberg’s classic result.
@@ 4.4.1 New peeling violation from ${\cal O}(G^{3})$ contribution
The starting point for the evaluation of the Fourier transform of the non-analytic part of the metric in Eq. (27) is Eq. (3.1). Due to the logarithmic branch cut, [^26] We recall that, as discussed in Sec. 2.2 and in footnote 12, logarithmic contributions are both related to the tail effect and to the contribution of Coulomb modes to the final-state graviton one-point function, Eq. (14). the residues overlap with a branch point. Thus, instead of Fig. 3, the contour for the $\omega$ integral is deformed as shown in Fig. 4, leading to
@@
where the ellipsis stands for terms of higher orders in $\rho\ll{|\vec{x}|^{-1}}\ll\text{(any other scale)}$.
We note that $\log\rho$ cancels out between the pole and the discontinuity, Eqs. (LABEL:eq:IR_residue) and (4.4.1), respectively. Moreover, the explicit $t$ dependence can be recast as a dependence on $|\vec{x}|$ by writing
@@
Adding up the two contours in Fig 4 and Eq. (105), it follows that, at ${\cal O}(G^{3})$ order, the terms proportional to $\ln(|\vec{x}|)\,{\widetilde{h}}^{\mu\nu}_{\textrm{tree}}(k)$ can be absorbed into a redefinition of the retarded time coordinate,
@@
The second line of Eq. (4.4.1) is governed by the leading off-shell properties of ${\widetilde{h}}_{\textrm{tree}}^{\mu\nu}(k)$, i.e. the part of the graviton one-point function in which momenta are treated as off-shell while not being proportional to powers of $k^{2}$ or of the type $k^{(\mu}V^{\nu)}$ with some vector $V$. The latter is projected out when carrying out the analysis above directly for the Weyl tensor of for the NP scalars while the latter, as discussed, can contribute only subleading terms in the large $|\vec{x}|$ expansion.
@@
$\Psi_{0}^{\not\perp}(x)=\frac{G}{|\vec{x}|}\int\frac{d\mathsf{k}}{2\pi}\mathsf{k}^{d-3}\int dn_{\perp}\int\frac{dy}{2\pi}e^{-\mathsf{k}y|\vec{x}|}$ (113)
@@
$L\cdot k=\frac{\omega}{\sqrt{2}}(1-z)\qquad\varepsilon_{+}\cdot k=-\frac{\omega}{\sqrt{2}}e^{-\phi}\sqrt{1-z^{2}}\ .$ (115)
The three terms in Eq. (113) have different $\phi$ dependence, and the integral over $\phi$ vanishes unless each factor of $e^{i\phi}$ pulls its conjugate out of $\partial_{\omega}((p_{1}+p_{2})\cdot k\,{\widetilde{h}}^{\mu\nu}_{\text{tree}}(k))$. As in the previous section, the structure of spherical harmonics implies that each such factor comes with at last one factor of $\sqrt{1-z^{2}}$. Overall, each of the three terms starts out as $(1-z)^{2}\sim y^{2}$ in a small-$y$ expansion and thus yields a factor of $|\vec{x}|^{-3}$. Together with the overall factor in Eq. (113) it follows that
@@
Thus, we find that the large-distance scaling of the radiation contribution to $\Psi_{0}$ at ${\cal O}(G^{3})$ departs from Sachs’s peeling property. This departure is, however, for a different reason than the Coulomb region. [^28] We also recall that, as we discussed in Sec. 4.2.2, the Coulomb-region contributions are confined to the tree level metric while the contribution found here is at a higher order. This departure is due to the non-analyticity of the spacetime metric from the long-range nature of gravitational interactions, which bring off-shell contributions of the graviton one-point function in the large-$|\vec{x}|$ expansion. [^29] We note that, since it is entirely due to outgoing gravitons, the same phenomena should also occur in the bound case. See Ref. Ivanov:2025ozg for a discussion of tails in the bound and unbound context.
@@
$\Psi_{1}^{\not\perp}(x)$ $=\frac{G}{|\vec{x}|}\int\frac{d\mathsf{k}}{2\pi}\mathsf{k}^{d-3}\int dn_{\perp}\int\frac{dy}{2\pi}e^{-\mathsf{k}y|\vec{x}|}\frac{\partial}{\partial\omega}\Big[(p_{1}+p_{2})\cdot k\,\big((L\cdot k)^{2}\varepsilon_{+}^{\mu}N^{\nu}$ (117)
@@
Using Eq. (115), $N\cdot k=\frac{\omega}{\sqrt{2}}(1+z)$ and the properties of spherical harmonics, it is not difficult to see that the first term on the second line of Eq. (117) and the image of the second term under the map $(\varepsilon_{+}\leftrightarrow N)$ imply that at ${\cal O}(G^{3})$ scales as
@@
This constitutes a stronger violation of the peeling property than that found in the Coulomb region which, as previously discussed, is confined to $\Psi_{0}$ and at ${\cal O}(G^{2})$. For the same reason, the Coulomb-region contribution cannot cancel this. Eq. (118) is one of our main technical results.
An analogous calculation reveals that $\Psi_{2}$ falls off as $|\vec{x}|^{3}$, thus obeying Sachs’s peeling property. The analysis in Sec. 4.3 regarding the effects of possible corrections to the tetrad. Since the corrections to the tetrad are expected to be ${\cal O}(G/|\vec{x}|)$ and ${\cal O}(G^{2}/|\vec{x}|)$ (which is the order of the correction to the metric itself), they will mix with the leading and next-to-leading terms in the NP scalars. Such terms do preserve peeling. Thus, the argument presented in Sec. 4.3.4 applies identically here.
The Weyl tensor is a non-linear function of the metric fluctuations. For example, at quadratic order in $h_{\mu\nu}=g_{\mu\nu}-\eta_{\mu\nu}$ it is given by
@@
The NP scalars pick up the corresponding projections of these terms on the null tetrad. One may wonder if such contributions can alter the conclusion that the non-analytic contributions to the metric are a novel source of peeling violation at ${\cal O}(G^{3})$. We will argue here that this cannot be the case.
In the classical limit expectation values of products of operators factorize into products of expectation values, to these contributions can be computed in terms of the classical metric fluctuations. Since we are estimating the bilinear contributions to the linearized Weyl tensor at a fixed order in $G$, the two metric fluctuations must be of lower order. At ${\cal O}(G^{3})$ they can only be either the expanded Schwarzschild metric or the tree-level scattering metric. Each factor has an ${\cal O}(1/|\vec{x}|)$ falloff, see e.g. Sec. 4.3.1. Direct calculations (or dimensional analysis) indicate that each derivative of the (expanded) Schwarzschild metric is suppressed by ${\cal O}(1/|\vec{x}|)$ relative to the metric. Recalling the form of the tree-level scattering metric as a Fourier transform, it follows that its derivative (or the corresponding momentum) points along a null direction; at subleading orders there is a departure from the light-like direction at least of ${\cal O}(1/|\vec{x}|)$. Thus, the two-fluctuation part of the Weyl tensor falls off at least like ${\cal O}(1/|\vec{x}|^{3})$ because in (119) at least one derivative acts on a Schwarzschild metric and the other derivative points in a null direction up to ${\cal O}(1/|\vec{x}|)$ corrections.
To complete the argument we must inspect the effects of the projection of the Weyl tensor onto the null tetrad. The four free indices of the Weyl tensor in (119) can be supplied either by the background Minkowski metric or, as dictated by soft-region power counting, by a pair of graviton momenta. The null nature of the tetrad projects out the contributions bilinear in the Minkowski metric for both $\Psi_{0}$ and $\Psi_{1}$. In the remaining terms, two tetrad vectors are contracted with derivatives on a metric fluctuation – either of the expanded Schwarzschild metric or of the tree-level scattering metric. Since both of them are at an order in $G$ at which the peeling property is observed, we can infer that the scalar product of $L$ and $\varepsilon_{+}$ vectors and a derivative falls off at least as ${\cal O}(1/|\vec{x}|)$.  [^30] We already discussed there properties explicitly in Eq. (115) and below Eq. (117). Since for $\Psi_{0}$ both derivatives are contracted with either $L$ and $\varepsilon_{+}$ and for $\Psi_{1}$ at most one derivative is contracted with $N$, see Eqs. (1), we conclude that the bilinear corrections to the NP scalars fall off as
$\Psi_{0}\Big|_{\text{bilinear}}\sim{\cal O}(1/|\vec{x}|^{5})\qquad\Psi_{0}\Big|_{\text{bilinear}}\sim{\cal O}(1/|\vec{x}|^{4})\ .$ (120)
@@ 4.4.2 Fourier transform with exponentiated logarithms
Since Eq. (121) is nonanalytic in $k^{2}$, the calculation follows the same steps as in Sec. 4.4.1. It is not difficult to notice that the phase in Eq. (121) regulates the integration along the discontinuity of the $\omega$ integral in Eq. (106). This implies that there is no need to consider separately the contours wrapping the residues and the branch cuts because the former can be discarded as we shrink the radius of the contour to zero. Thus, the only contribution to the Fourier transform comes from the branch cut:
@@
This expression shows clearly how the exponentiation of the IR divergences implies the exponentiation of the $\ln\mu|\vec{x}|$ contributions (upon writing the time in terms of the retarded time and expanding in $u/|\vec{x}|$). [^31] To write the entire prefactor as a exponential, one may use the identity: $\frac{1}{\pi}\sin(\pi a)\Gamma[a]=e^{-\gamma_{E}a-\sum_{n=2}^{\infty}\frac{\zeta(n)}{n}a^{n}}\ ,$ (126) where $\zeta(n)$ is the Riemann zeta function. It also indicates that

### v2

@@ 4.3.3 The NP scalars: analytic contributions
In particular, the integral over the azimuthal angle vanishes unless the integrand is independent of $\phi$. Since the angle $\phi$ enters [^28] A way to understand this dependence is from the structure of spherical harmonics. only via the complex combination of coordinates $x\pm iy=e^{\pm i\phi}\sqrt{1-z^{2}}$ and, as discussed, the coefficient functions $\alpha_{i}$ are regular at $z=1$, the expansion of the NP scalar $\Psi_{4}$ around $z=1$ is
@@
This concludes the main part of our proof that the part of the momentum-space metric that is analytic at $k^{2}=0$ yields a space-time metric that obeys Sachs’s peeling property.
@@
Each of the coefficients $B$ and $C$ is at least $\sim G/|\vec{x}|$ because the first correction to the metric, corresponding to the direct sum of two noninteracting Schwarzschild black holes, is of this order. Moreover, we can use the leading values of the other two coefficients, $A=D=1$. Thus, while the falloff of $\Psi_{i\geq 1}$ is slower than the expected falloff of $\Psi_{0}$, the fall-off of the coefficients $B$ and $C$ makes up for the difference. Consequently, the conclusion of Sec. 4.3.3 that the analytic contributions to the metric exhibit the peeling property are unaffected by redefinitions of the tetrad. We expect that this holds beyond the leading order in the deformation, so that such mixing of NP scalars due to the nontriviallity of the nullness and orthogonality constraints on the tetrad does not, on its own, lead to departures from the peeling property of the Weyl tensor.
In this section, we focused on the part of the momentum-space metric that is analytic in $k^{2}$. Thus, the only singularity in this part of the metric originates in the outgoing-graviton propagator. Together with the second essential ingredient, the $z\rightarrow 1$ behavior of the coefficient functions $\alpha_{i}$ in Eq. (45), it led to the position-space metric obeying Sachs’s peeling property. In the next section, we will discuss nonanalytic features of the final-state graviton one-point function and their consequences on the large-distance fall-off of the NP scalars.
@@ 4.4 Radiation region: non-analytic contributions and new peeling violation
We recall that the non-analytic terms can be constructed via the mapping identified in Sec. 2.2 at ${\cal O}(G^{3})$, see Eq. (26). In general, this mapping does not fully reconstruct the momentum-space metric. Indeed, the on-shell one-point function misses terms proportional to integer powers of the squared graviton momentum, $k^{2}$, and to $\varepsilon_{\mu\nu}(k)k^{\nu}$. The tensor factor in Eqs. (21) and (22) remedies the situation, producing a result proportional to $k^{2}$ when contracted with the graviton momentum. Moreover, any terms linear in $k^{\mu}$ that survive after this contraction drop out when evaluating the Weyl tensor and the NP scalars. Last, upon integration, these terms may contain non-analytic parts, but the $k^{2}$ factor brings an extra suppression in powers of $1/|\vec{x}|$ with respect to the non-analytic terms that arise from the on-shell mapping.
@@ 4.4.1 New peeling violation from ${\cal O}(G^{3})$ contribution
The starting point for the evaluation of the Fourier transform of the non-analytic part of the metric in Eq. (26) is Eq. (3.1). Due to the logarithmic branch cut, [^30] We recall that, as discussed in Sec. 2.2 and in footnote 15, logarithmic contributions to the final-state graviton one-point function, Eq. (13), are both related to the tail effect and to the contribution of BMS modes at $\mathcal{I}^{-}$ (which in turn correspond to different configurations at finite time). the residues overlap with a branch point. Thus, instead of Fig. 3, the contour for the $\omega$ integral is deformed as shown in Fig. 4, leading to
@@
where the ellipsis stands for terms of higher orders in $\rho\ll{|\vec{x}|^{-1}}\ll\text{(any other scale)}$. We note that $\log\rho$ cancels out between the pole and the discontinuity, Eqs. (LABEL:eq:IR_residue) and (4.4.1), respectively. Moreover, the explicit $t$ dependence can be recast as a dependence on $|\vec{x}|$ by writing
@@
Adding up the two contours in Fig. 4 and Eq. (103), it follows that, at ${\cal O}(G^{3})$ order, the terms proportional to $\ln(|\vec{x}|)\,{\widetilde{h}}^{\mu\nu}_{\textrm{tree}}(k)$ can be absorbed into a redefinition of the retarded time coordinate,
@@
The second line of Eq. (4.4.1) is governed by the leading off-shell properties of ${\widetilde{h}}_{\textrm{tree}}^{\mu\nu}(k)$, i.e. the part of the graviton one-point function in which momenta are treated as off-shell while not being proportional to powers of $k^{2}$ or of the type $k^{(\mu}V^{\nu)}$ with some vector $V$. The latter is projected out when carrying out the analysis above directly for the Weyl tensor of for the NP scalars, while the former can contribute only subleading terms in the large-$|\vec{x}|$ expansion.
@@
$\Psi_{0}^{\not\perp}(x)=\frac{G}{|\vec{x}|}\int\frac{d\mathsf{k}}{2\pi}\mathsf{k}^{d-3}e^{-i\mathsf{k}u}\int dn_{\perp}\int\frac{dy}{2\pi}e^{-\mathsf{k}y|\vec{x}|}$ (111)
@@
$L\cdot k=\frac{\omega}{\sqrt{2}}(1-z)\qquad\varepsilon_{+}\cdot k=-\frac{\omega}{\sqrt{2}}e^{-i\phi}\sqrt{1-z^{2}}\ .$ (113)
The three terms in Eq. (111) have different $\phi$ dependence, and the integral over $\phi$ vanishes unless each factor of $e^{i\phi}$ pulls its conjugate out of $\partial_{\omega}((p_{1}+p_{2})\cdot k\,{\widetilde{h}}^{\mu\nu}_{\text{tree}}(k))$. As in the previous section, such term comes with at last one factor of $\sqrt{1-z^{2}}$. Overall, each of the three terms starts out as $(1-z)^{2}\sim y^{2}$ in a small-$y$ expansion and thus yields a factor of $|\vec{x}|^{-3}$. Together with the overall factor in Eq. (111) it follows that
@@
Thus, we find that the large-distance scaling of the radiation contribution to $\Psi_{0}$ at ${\cal O}(G^{3})$ departs from Sachs’s peeling property. This departure is, however, for a different reason than the Coulomb region. [^32] We also recall that, as we discussed in Sec. 4.2.2, the Coulomb-region contributions are confined to the tree-level metric while the contribution found here is at a higher order. This departure is due to the non-analyticity of the spacetime metric from the long-range nature of gravitational interactions, which bring off-shell contributions of the graviton one-point function in the large-$|\vec{x}|$ expansion. [^33] We note that, since it is entirely due to outgoing gravitons, the same phenomena should also occur in the bound case. See Ref. Ivanov:2025ozg for a discussion of tails in the bound and unbound context. On the other hand, the same strategy cannot be straightforwardly applied to the bound case, as it is not possible to define an asymptotic bound state for macroscopic gravitational systems (strictly speaking, bound states in relativistic classical mechanics do not exist due to radiation emission).
@@
$\Psi_{1}^{\not\perp}(x)$ $=\frac{G}{|\vec{x}|}\int\frac{d\mathsf{k}}{2\pi}\mathsf{k}^{d-3}e^{-iu\mathsf{k}}\int dn_{\perp}\int\frac{dy}{2\pi}e^{-\mathsf{k}y|\vec{x}|}\frac{\partial}{\partial\omega}\Big[(p_{1}+p_{2})\cdot k\,\big((L\cdot k)^{2}\varepsilon_{+}^{\mu}N^{\nu}$ (115)
@@
Using Eq. (113), $N\cdot k=\frac{\omega}{\sqrt{2}}(1+z)$ and the properties of spherical harmonics (see footnote 28 and the discussion around Eq. (98)), it is not difficult to see that the first term on the second line of Eq. (115) and the image of the second term under the map $(\varepsilon_{+}\leftrightarrow N)$ imply that at ${\cal O}(G^{3})$ scales as
@@
This constitutes a stronger violation of the peeling property than that found in the Coulomb region. For the same reason, the Coulomb-region contribution cannot cancel this. Eq. (116) is one of our main technical results.
The scaling of $\Psi_{0}$ and $\Psi_{1}$ at large $|\vec{x}|$ were given by simple dimensional arguments and the inspection of the angular dependence of the momentum-space Weyl tensor. These arguments are not sufficient to fix the scaling of $\Psi_{2}$, which is sensitive to terms $\propto k^{2}$ in the metric. An explicit computation of these contributions is needed to understand the contribution of the tails to the falloff of this NP scalar.
As this point, we should carefully provide an analysis similar to the one in Sec. 4.3 regarding the effects of possible corrections to the tetrad. Since the corrections to the tetrad are expected to be ${\cal O}(G/|\vec{x}|)$ and ${\cal O}(G^{2}/|\vec{x}|)$ (which is the order of the correction to the metric itself), they will mix with the leading and next-to-leading terms in the NP scalars. Such terms preserve peeling. Thus, the argument presented in Sec. 4.3.4 applies identically here. We also analyze the possible contribution from the non-linearities of the Weyl tensor on the metric. Indeed, at quadratic order in $h_{\mu\nu}=g_{\mu\nu}-\eta_{\mu\nu}$ the Weyl tensor is given by
@@
The NP scalars pick up the corresponding projections of these terms on the null tetrad. One may wonder if such contributions can alter the conclusion that non-analytic contributions to the metric are a novel source of peeling violation at ${\cal O}(G^{3})$. We will argue here that this cannot be the case.
Since we are estimating the bilinear contributions to the linearized Weyl tensor at a fixed order in $G$, the two metric fluctuations must be of lower order. At ${\cal O}(G^{3})$ they can only be either the expanded Schwarzschild metric or the tree-level scattering metric. Each factor has an ${\cal O}(1/|\vec{x}|)$ falloff, see e.g. Sec. 4.3.1. Direct calculations (or dimensional analysis) indicate that each derivative of the (expanded) Schwarzschild metric is suppressed by ${\cal O}(1/|\vec{x}|)$ relative to the metric. Recalling the form of the tree-level scattering metric as a Fourier transform, it follows that its derivative (or the corresponding momentum) points along a null direction; at subleading orders there is a departure from the light-like direction at least of ${\cal O}(1/|\vec{x}|)$. Thus, the two-fluctuation part of the Weyl tensor falls off at least like ${\cal O}(1/|\vec{x}|^{3})$ because in (117) at least one derivative acts on a Schwarzschild metric and the other derivative points in a null direction up to ${\cal O}(1/|\vec{x}|)$ corrections. To complete the argument, we must inspect the effects of the projection of the Weyl tensor onto the null tetrad. The four free indices of the Weyl tensor in (117) can be supplied either by the background Minkowski metric or, as dictated by soft-region power counting, by a pair of graviton momenta. The null nature of the tetrad projects out the contributions bilinear in the Minkowski metric for both $\Psi_{0}$ and $\Psi_{1}$. In the remaining terms, two tetrad vectors are contracted with derivatives on a metric fluctuation – either of the expanded Schwarzschild metric or of the tree-level scattering metric. Since both of them are at an order in $G$ at which the peeling property is observed, we can infer that the scalar product of $L$ and $\varepsilon_{+}$ vectors and a derivative falls off at least as ${\cal O}(1/|\vec{x}|)$. [^34] We already discussed their properties explicitly in Eq. (113) and below Eq. (115). Since for $\Psi_{0}$ both derivatives are contracted with either $L$ and $\varepsilon_{+}$, and for $\Psi_{1}$ at most one derivative is contracted with $N$, see Eqs. (1), we conclude that the bilinear corrections to the NP scalars fall off as
$\Psi_{0}\Big|_{\text{bilinear}}\sim{\cal O}(1/|\vec{x}|^{5})\qquad\Psi_{1}\Big|_{\text{bilinear}}\sim{\cal O}(1/|\vec{x}|^{4})\ .$ (118)
@@ 4.4.2 Fourier transform with exponentiated logarithms
Since Eq. (119) is non-analytic in $k^{2}=0$, the calculation follows the same steps as in Sec. 4.4.1. It is not difficult to notice that the phase in Eq. (119) regulates the integration along the discontinuity of the $\omega$ integral in Eq. (104). This implies that there is no need to separately consider the contours wrapping the residues and the branch cuts because the former can be discarded as we shrink the radius of the contour to zero. Thus, the only contribution to the Fourier transform comes from the branch cut:
@@
This expression clearly shows how the exponentiation of the IR divergences implies the exponentiation of the $\ln\mu|\vec{x}|$ contributions (upon writing the time in terms of the retarded time and expanding in $u/|\vec{x}|$). [^35] To write the entire prefactor as a exponential, one may use the identity: $\frac{1}{\pi}\sin(\pi a)\Gamma[a]=e^{-\gamma_{E}a-\sum_{n=2}^{\infty}\frac{\zeta(n)}{n}a^{n}}\ ,$ (124) where $\zeta(n)$ is the Riemann zeta function. It also indicates that